# FM2DS: Few-Shot Multimodal Multihop Data Synthesis with Knowledge Distillation for Question Answering

### v2

@@ 5 Experiments and Results
This section evaluates our synthesized dataset against human-annotated datasets. In all experiments, the model was provided one in-context example during data synthesis, unless stated otherwise. We used GPT-4o in the synthesis pipeline. Models were evaluated using Exact Match (EM) for answer accuracy and F1 score for precision-recall balance in partial matches. See Appendix B for additional experimental details.
5.1 Comparison with Human-Annotated Datasets
Unlike previous dataset creation methods like MultiModalQA [35] and WebQA [5], our approach is fully automated, with no human intervention in data generation. This section provides a comparative analysis of the data quality synthesized by our method against these prior human-annotated datasets. We trained three models—LLaVA-1.6 [28, 27, 29], InternVL-2 [11, 12], and Idefics-2 [23, 25]—with varying sizes of the WebQA and MultiModalQA datasets and evaluated them on their respective test sets. Table 2 shows that samples generated by FM2DS help us train models that can handle MMQA better, even though its training samples include long documents. In contrast, training samples from MultiModalQA and WebQA consist of short, focused information snippets. We also notice that WebQA might be an easier task than MultiModalQA, as models tend to perform better on WebQA with fewer samples, even though it has a larger training set. Across all models and experiments with different numbers of training samples (where models trained on synthetic data were exposed to an equal or fewer number of real samples), the average improvement observed in EM was 1.81 in MultiModalQA and 1.96 in WebQA. While improvements in EM generally lead to increases in F1 score, there are cases where F1 decreases even as EM improves. This discrepancy may result from hallucinations in incorrect predictions, which reduce string overlap and, in turn, lower the F1 score.
Moreover, models achieve the same performance with less synthetic samples generated by FM2DS as those trained on the full training dataset, demonstrating faster convergence with fewer samples. We explore this in details in Section 5.2. Also note that larger models achieves better performance on the same number of synthetic samples, see the difference between Llava-1.6-13B and Llava-1.6-7B. GPT-4o achieves better results among large VLMs, likely because it was also used for data generation. For qualitative analysis, refer to Appendix F.
@@ 5.1 Comparison with Human-Annotated Datasets
Table 2: The result of comparing synthetic data with a human-annotated dataset across different models. For the smaller versions of the models, we tested with 5k samples, 10k samples, and the full training set (23.8k for MultiModalQA and 34.2k for WebQA) and for bigger models we tested on full training set and 10k samples. When comparing the model trained on the full dataset with synthetic data, the mentioned number of synthetic samples is the smallest (divisible by 1000) that outperforms the full dataset model. For real samples, we used the WebQA training set for testing on the WebQA test set, and similarly for MultiModalQA. The results show that models trained on synthetic data achieve higher EM across all experiments. Models with “None” are the default pretrained models. For additional results, see Appendix C.
@@ 5.2 Learning Efficiency Comparison
Figure 3: Comparison of learning from Synthesis Data and Real Data across 1k to 10k samples for InternVL-2-8B. As shown, with an equal number of samples, synthetic data generated by FM2DS outperforms real (human-annotated) samples in terms of both EM and F1 scores. Additionally, the gap between the two datasets narrows as the number of samples increases.
Figure 4: Comparison of learning from synthetic Data and Real Data across 1k to 10k samples for InternVL-2 8B. In this experiment, to make the real samples similar to ours, we used the full Wikipedia pages from Multimodal QA dataset.
To assess the comparative learning efficiency between synthesized and human-annotated datasets, we conducted systematic experiments using InternVL-2-8B across incremental training sample sizes ranging from 1k to 10k instances. For synthesized data, we used the same training data for both MultiModalQA and WebQA. For the human-collected datasets, we trained the model on each corresponding training set.
As illustrated in Figure 3, our synthesized dataset shows a substantial performance over real data at smaller sample sizes. However, as the number of real samples increases, this gap gradually narrows, though the synthesized dataset still achieves superior results overall. Notably, as the number of samples approaches 10k, the learning efficiency of the model trained on synthesized data declines more noticeably compared to real data. This slowdown may be attributed to the broader and more diverse distribution of general knowledge present in the synthesized data. While this diversity is beneficial for earlier stages of training, it can eventually lead to diminishing returns as the model begins to saturate. In contrast, real data, which often contains more focused and specific information, may continue to provide new, relevant patterns for the model to learn, maintaining a steadier learning pace [17, 31].
In a related experiment with MultiModalQA, we used full Wikipedia pages as training data by leveraging the dataset’s linked articles. Unfortunately, this setup could not be replicated with WebQA, as its source links predominantly reference WikiMedia pages without sufficient textual information. Results in Figure 4 show that, initially, the model trained on full Wikipedia pages for real samples performs better (in terms of improvement per 1k samples); however, after approximately 3k samples, this gap decreases. This indicates that, unlike our synthesized dataset, full-page real data in MultiModalQA lacks the generality offered by synthesized data. Synthesized data not only improves model learning by including queries but also maintains consistent learning speed, likely due to the diversity and quality of the samples.
5.3 Cross-Dataset Evaluation
To evaluate the generalizability and robustness of our synthesized data, we conducted a cross-dataset evaluation using the InternVL-2-8B model. For this experiment, we used our M2QA-Bench with 500 samples and matched it with 500 randomly selected samples from the MultiModalQA test set, including full Wikipedia pages for both. With evaluation sets now having a similar structure, we trained the model separately on 5k samples from our synthesized dataset and 5k training samples from the MultiModalQA dataset, both with full Wikipedia pages as context. As shown in Table 3, the model trained on our synthesized data outperformed the model trained on MultiModalQA samples across both test sets. This result highlights the broader generalizability of our dataset. Additionally, our benchmark proves to be more complex and diverse than MultiModalQA which has predominantly template-based questions.
@@ 5.3 Cross-Dataset Evaluation
Table 3: Cross-Dataset Evaluation Results With InternVL-2-8B for MultiModalQA and Our Synthesized Benchmark. Our generated benchmark is challenging, and the synthetic data produced by FM2DS is diverse and general, leading to improved cross-dataset performance.
6 Human Evaluation of Answer Validation
To evaluate the accuracy and impact of the automatic answer validation component of our pipeline, we conducted a human evaluation study. We randomly selected 100 samples and for each sample, we generated an answer using two methods: FM2DS and FM2DS w/o answer validation. We then invited 10 participants to contribute to our evaluation platform (see Appendix D) to determine which method produced the correct answer. Participants were instructed to verify which answer was correct by reviewing the associated Wikipedia pages. Once they determined the accuracy of each answer, they were asked to select one of four options: (1) Method 1’s answer is correct, (2) Method 2’s answer is correct, (3) both methods generated the correct answer, or (4) neither answer is correct. The results of the study are presented in Figure 5, shows a clear trend that answer validation improves the likelihood of correct answers.
7 Ablation Studies
7.1 Effect of Key Steps on Model Performance
To analyze the impact of various data filtering and validation steps on model performance, we evaluated the model across different configurations. Table 4 demonstrate the importance of each data refinement step on model performance. Question validation improved relevance by filtering questions tailored to the complex, multimodal, multihop requirements of the test sets, reducing hallucinations and boosting F1 by keeping answers concise and accurate. Answer validation and query generation proved crucial as answer validation filtered out incorrect samples, and queries supported learning by distilling knowledge from larger models, leading to improved EM and F1 scores. Finally, query validation reinforced training consistency by ensuring that queries followed the correct structure, minimizing confusion and helping the model accurately interpret questions. Together, these steps yielded the highest scores, underlining the cumulative impact of careful dataset curation.
@@ 7.1 Effect of Key Steps on Model Performance
QV 64.61 70.12 70.28 78.44
QV + AV 67.96 73.56 74.59 79.67
QV + AV + Qu 70.83 75.84 77.49 81.86
QV + AV + Qu + QuV 73.92 84.27 81.27 90.21
Table 4: Results of the FM2DS with and without key steps like query generation and verification. Document grouping, few-shot sample selection, question generation, and answer generation are included in all results. QV denotes Question Validation, AV Answer Validation, Qu Query, and QuV Query Validation.
7.2 The Effect of the Number of In-Context Documents
Table 5 presents the results of evaluating the Intervl-2 8B model with varying numbers of in-context documents on the MultiModalQA and WebQA datasets. In the zero-shot setting, FM2DS exhibits limited understanding of multimodal multi-hop question answering, and occasionally circumvents the validation step by simply generating a question that is not multihop. For example, ”Looking at the image of the Eiffel Tower, what engineering innovation allowed it to surpass previous structures in height?” prompts the model to use the image, but the answer is available in the page’s text on tall structures. As we move from zero-shot to one-shot, there is a significant boost in EM and F1 scores, reflecting improved performance with minimal context. The improvement from one-shot to two-shot is marginal, suggesting diminishing returns. With three in-context samples, the gains become minimal, indicating that additional samples beyond two provide little benefit. This diminishing return may stem from the model’s limited context window, which restricts its ability to fully utilize large in-context samples [22, 3].
Number of Shots MultiModalQA WebQA
EM F1 EM F1
zero-shot 66.42 75.1 71.05 79.3
one-shot 73.92 84.27 81.27 90.21
two-shot 74.3 83.71 81.5 91.37
three-shot 74.45 85.39 81.6 91.21
Table 5: Effect of the number of in-context documents on the performance of Intervl-2 8B on MultiModalQA and WebQA datasets.
7.3 Comparing Methods For Data Synthesis
To evaluate the effectiveness of different methods for synthetic data generation, we compared three prominent language models: GPT-4o, Claude 3.5 Sonnet, and Llama-3.2-90B, as shown in Table 6. Using Intervl-2-8B with 5K fine-tuning samples as our baseline model, we tested the quality of generated data on two distinct datasets: MultiModalQA and WebQA. The results, measured using EM and F1 scores, demonstrate that GPT-4o consistently outperforms other models across both datasets. We also see that Llama-3.2-90B shows competitive performance as an open-source model with less number of parameters, particularly in WebQA tasks. Claude 3.5 Sonnet generally yields lower scores across both datasets.
Model MultiModalQA WebQA
EM F1 EM F1
GPT-4o 73.92 84.27 81.27 90.21
Claude-3.5-Sonnet 68.35 76.01 76.98 83.86
Llama-3.2-90B 70.64 76.32 79.81 86.75
@@ 8 Conclusion & Future Works
We have presented a novel methodology for synthesizing high-quality data for multimodal multihop question answering, called FM2DS. Our approach addresses the limitations of existing methods, which focus on single-hop question answering with a single modality, and provides a robust framework for generating complex questions and answers that require reasoning over multiple sources of information with minimal human effort. We have demonstrated the effectiveness of our approach in creating a high-quality dataset that enables training models for multimodal multihop question answering. Our experiments have shown that the synthesized dataset allows models to achieve higher performance than when trained on real data collected by humans, with significant improvements in test scores. These results highlight the potential of our approach to improve the state-of-the-art in multimodal multihop question answering . Furthermore, our methodology provides a scalable and efficient solution for generating high-quality data, which is essential for training large language models. For future work, we plan to synthesize MMQA samples using sources beyond Wikipedia, including multilingual content, code snippets, videos, and other diverse information type.
@@ Appendix A Prompts
Prompt 1: The prompt for question generation defines what constitutes a multi-hop question and instructs the model to create a multimodal and multihop question based on the provided documents. It emphasizes that the question should require information from multiple modalities and documents to be answered, similar to given example(s).
@@ A.3 Query Generation
As illustrated in Prompt 3, in query generation, the model is tasked with explaining the step-by-step process used to extract relevant information from the documents and determine the answer based on the extracted snippets. This task emphasizes transparency by requiring the model to identify the relevant sections of each document and describe how information from multiple sources is combined to arrive at the correct answer, promoting accountability in the model’s reasoning process.
@@ Appendix B Experimental Settings
In this work, we conducted experiments on a cluster of 8 NVIDIA H100 80GB GPUs. The distributed setup allowed us to efficiently scale our fine-tuning process across multiple devices. The fine-tuning process was carried out using low-rank adaptation (LoRA) [19], a technique for efficient adaptation of pretrained models with low-rank matrices, reducing the number of trainable parameters. The key hyperparameters used in the fine-tuning procedure include a learning rate of 1e-4, a batch size of 8 per device (totaling 64 across 8 devices), LoRA rank set to 8, LoRA alpha set to 32, a weight decay of 0.01, and the number of epochs was 5. Additionally, AdamW optimizer was used with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=1e-8$. The models were fine-tuned using mixed-precision training to take full advantage of the 80GB memory on each H100 GPU. For inference time, we set the temperature to 0.7, which strikes a balance between randomness and coherence in the model’s responses, producing more varied outputs without sacrificing too much quality. This setup ensured efficient usage of computational resources while maintaining high model performance.
Appendix C Investigating FM2DS on Other Models
In addition to the models discussed in Section 5, we explored other model families, including Idefics3 [24], mPLUG-DocOwl-1.5 [18], and Phi-3.5-Vision-Instruct [1], as well as larger versions within the explored families presented in Table 2. The results in Table 7 demonstrate the reliability of our data synthesis approach, which consistently enhances model performance across all models and sizes compared to an equivalent number of real samples.

### v4

@@ 5 Experiments and Results
This section compares our synthesized dataset to human-annotated ones. All experiments used one in-context example during synthesis (see Appendix C for effects of varying number in-context examples). GPT-4o was used in the pipeline (Appendix D shows results with other LVLMs). Models were evaluated using Exact Match (EM) for accuracy and F1 for partial match quality. Further experimental details can be found in Appendix B.
@@
Table 3: Comparison of synthetic and human-annotated data across various models. Smaller models were evaluated with 5k, 10k, and full training sets (23.8k for MultiModalQA, 34.2k for WebQA), while larger models used 10k and full sets. For each model, the listed synthetic sample size is the smallest (divisible by 1k) that outperforms the same model trained on the full human-labeled set. For real samples, we used the WebQA training set for testing on the WebQA test set, and similarly for MultiModalQA. “None” indicates default pretrained models. For more results, see Appendix E.
5.1 Comparison with Human-Annotated Datasets
Unlike prior methods like MultiModalQA (Talmor et al., 2021) and WebQA (Chang et al., 2021), our approach is fully automated with no human involvement (minimal human evaluation was used in the M2QA-Bench only). This section compares the quality of our synthesized data against these human-annotated datasets. We trained LLaVA-1.6 (Liu et al., 2023b; a; 2024), InternVL-2 (Chen et al., 2023b; 2024c), and Idefics-2 (Laurençon et al., 2023; 2024b) on varying sizes of WebQA and MultiModalQA, evaluating on their respective test sets. We also trained the same models on FM2DS-generated training data and evaluated them on the same test sets to assess the effectiveness of the synthesized data in comparison to human-annotated datasets.
As shown in Table 3, models trained on FM2DS data outperform those trained on original datasets, despite using longer documents. WebQA seems easier than MultiModalQA, with better performance from fewer samples. On average, EM improved by 1.81 for MultiModalQA and 1.96 for WebQA using equal or fewer synthetic samples. While EM gains often lead to higher F1, some cases show F1 drops due to hallucinated answers reducing string overlap. Models trained on fewer synthetic samples from FM2DS match the performance of those trained on full datasets, showing faster convergence (see Section 5.2). Larger models also perform better with the same synthetic data—e.g., LLaVA-1.6-13B vs. 7B. GPT-4o leads among large LVLMs, likely due to its role in data generation (Refer to Appendix J for qualitative analysis). When comparing these human annotated data with our synthesized data, a common question is: ”Why not paraphrase existing datasets instead of synthesizing new ones?”. The answer is paraphrasing doesn’t enable domain adaptation, and as shown in Appendix K, fine-tuning on our synthetic data outperforms training on a paraphrased version of MultiModalQA.
Figure 3: (a) and (b): EM and F1 comparison on 1k–10k samples for InternVL-2-8B shows that FM2DS’s synthetic data outperforms human-annotated data, with the gap narrowing as sample size increases. (c): Similar comparison using full Wikipedia pages from the MultimodalQA dataset to match our synthetic data format.
@@ 5.2 Learning Efficiency Comparison
To evaluate learning efficiency, we ran experiments with InternVL-2-8B using incremental training sizes from 1k to 10k (in 1k steps) on both synthetic and human-annotated data. For real data, we used the same dataset for training and testing—e.g., when testing on WebQA, training samples were taken from the WebQA training set.
As shown in Figure 3(a & b), our synthesized data outperforms real data at smaller training sizes, though the gap narrows as sample size grows. Near 10k samples, learning efficiency with synthetic data declines more than with real data—likely due to its broader knowledge coverage. While this diversity aids early learning, it can lead to saturation, unlike real data, which offers more focused patterns and sustains steady learning (Hong et al., 2023; Maini et al., 2024).
In a related experiment on MultiModalQA, we used full Wikipedia pages via linked articles as training data instead of information snippets. This wasn’t possible for WebQA, as its source links mostly point to WikiMedia pages with limited text. Figure 3(c) shows that models trained on full Wikipedia pages initially achieve better improvement per 1k samples; however, this advantage declines after approximately 3k samples. This suggests that, while full-page real data offers some early benefits over short snippets, it still lacks the generality and consistency of our synthesized dataset. In contrast, the synthesized data—with its built-in queries and diverse content—continues to support steady learning, likely due to its broader coverage and higher quality.
@@
Table 4: Cross-Dataset Evaluation Results With InternVL-2-8B for MultiModalQA and Our Synthesized Benchmark.
5.3 Cross-Dataset Evaluation
To assess the generalizability of our synthesized data, we conducted a cross-dataset evaluation using InternVL-2-8B. We used 1k samples from M2QA-Bench and 1k from the MultiModalQA test set, both with full Wikipedia pages as sources. The model was trained separately on 5k samples from our dataset and 5k from MultiModalQA, using the same input format. As shown in Table 4, the model trained on our data outperformed the one trained on MultiModalQA across both test sets, demonstrating stronger generalization. This also reflects the greater complexity and diversity of our benchmark, compared to MultiModalQA’s template-based questions. See Appendix H to see the performance on model trained on FM2DS’s data when using on a out of domain MMQA dataset.
5.4 Key Stages in FM2DS
@@
Question Validation 64.61 70.12 70.28 78.44
+ Answer Validation 67.96 73.56 74.59 79.67
+ Query 70.83 75.84 77.49 81.86
+ Query Validation 73.92 84.27 81.27 90.21
Table 5: Results of the FM2DS with and without key steps like query generation and verification. All other steps are included in all of the results. The plus sign (”+”) at the start each steps means all the previous steps were included as well.
To assess the impact of each data filtering and validation step, we evaluated InternVL-2-8B on 5k training samples under different configurations. As shown in Table 5, each step contributed to performance gains. Question validation improved relevance by filtering questions tailored to the complex, multimodal, multihop requirements of the test sets, and reduced hallucinations, boosting F1 with more accurate answers. Answer validation removed incorrect samples, while query generation distilled knowledge from larger models, improving EM and F1. Query validation reinforced consistency by ensuring proper structure. Refer to Appendix L for statistics like ”Rejection Rate” on the validation stages used to remove specific samples during data synthesis.
5.5 M2QA-Bench Evaluation Results
Figure 4: Bar plot showing EM and F1 scores of multimodal models on M2QA-Bench.
To assess the complexity of M2QA-Bench and compare model performance, we evaluated a diverse set of models, as shown in Figure 4. GPT-4o stands out, outperforming even larger models like LLaMa-3.2-90B and Claude-3.5-Sonnet by a notable margin. Interestingly, smaller models in the 4B–8B range, particularly those from the Phi family (Abdin et al., 2024a; b), also achieve competitive results despite their scale. GPT-4o’s advantage may stem in part from its involvement in the initial data generation, but this remains an open question, as human annotators thoroughly reviewed and corrected the dataset to ensure fairness, reduce bias, and maintain high quality.
6 Human Evaluation of Answer Validation
To evaluate the accuracy and impact of our automatic answer validation component, we conducted a human evaluation study. After generating questions for 100 samples, we continued the pipeline using two methods: FM2DS with and without answer validation. These samples were divided into four batches of 25, with each batch evaluated by three participants to mitigate potential bias or human errors.
In total, 12 participants contributed through our evaluation platform (see Appendix F), assessing which method produced the correct answer.
Participants were instructed to verify the correctness of each answer by reviewing the associated Wikipedia pages. Once they determined the accuracy of each answer, they were asked to select one of four options: (1) Method 1’s answer is correct, (2) Method 2’s answer is correct, (3) both methods generated the correct answer, or (4) neither answer is correct (Method 1 and Method 2 were randomly assigned to the datasets generated with and without answer validation, respectively.). The results, shown in Figure 5, reveal a clear trend: answer validation increases the likelihood of correct answers, even though the model without validation still produced correct answers in over 60% of cases. The Fleiss’ Kappa was 0.91—indicating strong agreement—though this is expected, as the task involved factual questions with definitive answers rather than subjective judgments.
@@ 7 Conclusion & Future Works
We present FM2DS, a novel methodology for synthesizing high-quality data for multimodal multihop question answering. Unlike existing approaches that are limited to single-hop and single-modality settings, FM2DSgenerates complex QA pairs that require reasoning over multiple modalities and sources, with minimal human intervention. Our framework enables the creation of a large-scale dataset that significantly boosts model performance, surpassing models trained on human-curated data in terms of test accuracy. These results demonstrate the effectiveness of synthetic data for advancing the state of the art in multimodal multihop QA. Additionally, FM2DSoffers a scalable and efficient solution for training data-hungry language models. For future work, we plan to synthesize MMQA samples using sources beyond Wikipedia, including multilingual content, code snippets, videos, and other diverse information type.
Ethics Statement
We recruited human evaluators to support our research and compensated them at a rate of \$10 per hour. All evaluators provided informed consent prior to participation and were informed about the nature of the task. The annotation process adhered to ethical research guidelines, ensuring fair compensation and voluntary participation. No personally identifiable information was collected.
We also recruited three human annotators to verify and clean our M2QA-Bench data. These annotators were compensated at a rate of \$25 per hour for their expertise and time. All annotators provided informed consent and were briefed on the goals and procedures of the task. The annotation process followed ethical research guidelines, ensuring fair compensation, data confidentiality, and voluntary participation throughout.
@@ Appendix A Prompts
Prompt 1: The prompt for question generation defines what constitutes a multi-hop question and instructs the model to create a multimodal and multihop question based on the provided documents. It emphasizes that the question should require information from multiple modalities and multiple given documents to be answered, similar to the given example(s).
@@ A.3 Query Generation
As illustrated in Prompt 3, in query generation, the model is tasked with explaining the step-by-step process used to extract relevant information from the documents and determine the answer based on the extracted snippets. This task emphasizes transparency by requiring the model to identify the relevant sections of each document and describe how information from multiple sources is retrieved and combined to arrive at the correct answer, promoting explainability in the model’s reasoning process.
@@ Appendix B Experimental Settings
In this work, we conducted experiments on a cluster of 8 NVIDIA H100 80GB GPUs. The distributed setup allowed us to efficiently scale our fine-tuning process across multiple devices. The fine-tuning process was carried out using low-rank adaptation (LoRA) (Hu et al., 2022), a technique for efficient adaptation of pretrained models with low-rank matrices, reducing the number of trainable parameters. The key hyperparameters used in the fine-tuning procedure include a learning rate of 1e-4, a batch size of 8 per device (totaling 64 across 8 devices), LoRA rank set to 8, LoRA alpha set to 32, a weight decay of 0.01, and the number of epochs was 5. Additionally, AdamW optimizer was used with $\beta_{1}=0.9$, $\beta_{2}=0.98$, and $\epsilon=1e-8$. The models were fine-tuned using mixed-precision training to take full advantage of the 80GB memory on each H100 GPU. For inference time, we set the temperature to 0.7, which strikes a balance between randomness and coherence in the model’s responses, producing more varied outputs without sacrificing too much quality. This setup ensured efficient usage of computational resources while maintaining high model performance.
Appendix C The Effect of the Number of In-Context Examples
Number of Shots MultiModalQA WebQA
EM F1 EM F1
zero-shot 66.42 75.1 71.05 79.3
one-shot 73.92 84.27 81.27 90.21
two-shot 74.3 83.71 81.5 91.37
three-shot 74.45 85.39 81.6 91.21
Table 6: Effect of the number of in-context documents on the performance of Intervl-2-8B on MultiModalQA and WebQA datasets.
Table 6 presents the results of evaluating the Intervl-2-8B model with varying numbers of in-context examples on the MultiModalQA and WebQA datasets. In the zero-shot setting, FM2DS exhibits limited understanding of multimodal multi-hop question answering, and occasionally circumvents the validation step by simply generating a question that is not multihop. For example, ”Looking at the image of the Eiffel Tower, what engineering innovation allowed it to surpass previous structures in height?” prompts the model to use the image, but the answer is available in the page’s text on tall structures. As we move from zero-shot to one-shot, there is a significant boost in EM and F1 scores, reflecting improved performance with minimal context. The improvement from one-shot to two-shot is marginal, suggesting diminishing returns. With three in-context samples, the gains become minimal, indicating that additional samples beyond two provide little benefit. This diminishing return may stem from the model’s limited context window, which restricts its ability to fully utilize large in-context samples (Kaplan et al., 2020; Bertsch et al., 2024).
Appendix D Comparing Methods For Data Synthesis
To evaluate the effectiveness of different methods for synthetic data generation, we compared three prominent language models: GPT-4o, Claude 3.5 Sonnet, and Llama-3.2-90B, as shown in Table 7. Using Intervl-2-8B with 5K fine-tuning samples as our baseline model, we tested the quality of generated data on two distinct datasets: MultiModalQA and WebQA. The results, measured using EM and F1 scores, demonstrate that GPT-4o consistently outperforms other models across both datasets. We also see that Llama-3.2-90B shows competitive performance as an open-source model with less number of parameters, particularly in WebQA tasks. Claude 3.5 Sonnet generally yields lower scores across both datasets. As shown in Figure 4, Claude-3.5-Sonnet outperforms Llama-3.2-90B, which may be attributed to differences in the tasks that were included during their respective training phases. This observation can be further investigated.
Model MultiModalQA WebQA
EM F1 EM F1
GPT-4o 73.92 84.27 81.27 90.21
Claude-3.5-Sonnet 68.35 76.01 76.98 83.86
Llama-3.2-90B 70.64 76.32 79.81 86.75
Appendix E Investigating FM2DS on Other Models
In addition to the models discussed in Section 5, we explored other model families, including Idefics3 (Laurençon et al., 2024a), mPLUG-DocOwl-1.5 (Hu et al., 2024), and Phi-3.5-Vision-Instruct (Abdin et al., 2024a), as well as larger versions within the explored families presented in Table 3. The results in Table 8 demonstrate the reliability of our data synthesis approach, which consistently enhances model performance across all models and sizes compared to an equivalent number of real samples.